# Kalb-Ramond field, black holes and black strings in (2 + 1)D

### v4

@@ 3.3. Positive coupling
Fig. 4: The plot depicts the global structure of the black string solutions with $10<\gamma<$ 1. $\rho_{\pm}$ represent the horizons. The solutions have a time-like singularity at $\rho=\rho_{0}$. $\rho_{0}$ is represented by zigzag lines. For arbitrarily small positive $\gamma$, the singularity is at arbitrarily large $\rho=\rho_{0}$.
@@
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho_{+}^{2}-\rho^{2})(\rho_{-}^{2}-\rho^{2})}-{(\rho^{2}_{+}-\rho^{2}_{-})(\rho^{2}_{-}-\rho^{2})dt^{2}\over 4l^{2}\rho^{2}}+{(\rho_{+}^{2}-\rho_{-}^{2})\rho^{2}\over 4(\rho^{2}_{+}-\rho^{2})}\left(dx-{\rho_{-}\rho_{+}\over l\rho^{2}}dt\right)^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}={\rho_{+}^{2}-\rho_{-}^{2}\over 4l}={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,\quad e^{-2\phi}=4\left({\rho_{+}^{2}-\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right).}$ (3.33)3.33( 3.33 )
@@
$R={4\over l^{2}}\left({\rho_{+}^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{+}^{2}}\right).$ (3.34)3.34( 3.34 )
Thus, for an observer residing in the spacetime region $\rho<\rho_{-}$, i.e., the exterior region, the solution (3.33) has a singularity at $\rho=\rho_{0}=\rho_{+}$ beyond the horizon, i.e., $\rho_{0}>\rho_{-}$. Thus, in view of this observer, (3.33) describes a black hole. Since $x$ is compact, the solution contains closed time-like curves in the region $\rho>\rho_{0}=\rho_{+}$. Thus, they are located in the interior region. We note that the $B$ field, i.e. $B_{tx}$, is a non-zero constant. We can fix the $B$ field to zero using the gauge freedom. The Penrose diagram is given in Fig. 5.
Fig. 5: The plot depicts the global structure of the black string solution with $\gamma=1$. $\rho_{-}$ represents the horizon. The solution has a space-like singularity at $\rho_{0}=\rho_{+}$. $\rho_{0}$ is represented by zigzag lines.
@@ 3.4. New black hole solutions
$\eqalign{t&=\widehat{t}{\tenrm cosh}(\alpha)+l\widehat{x}{\tenrm sinh}(\alpha),\cr lx&=l\widehat{x}{\tenrm cosh}(\alpha)+\widehat{t}{\tenrm sinh}(\alpha),}$ (3.35)3.35( 3.35 )
@@
$\tanh(2\alpha)={2\rho_{+}\rho_{-}\over\rho_{+}^{2}+\rho_{-}^{2}}.$ (3.36)3.36( 3.36 )
@@
$\eqalign{d{\widehat{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-l^{-2}e^{2\phi}(\rho^{2}-\rho_{+}^{2})d{\widehat{t}}^{2}+e^{2\phi}(\rho^{2}-\rho_{-}^{2})d{\widehat{x}}^{2},\cr B&=B_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}}\wedge d{\widehat{x}},\cr B_{{\widehat{t}}{\widehat{x}}}&={1\over 2}e^{2\phi}\left({\rho_{+}^{2}-\rho_{-}^{2}\over l}\right)\left[\gamma-1-2\left({\rho^{2}-\rho_{+}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right)\right],\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\gamma)+\rho_{-}^{2}(1-\gamma)-2\rho^{2}\right],\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,}$ (3.37)3.37( 3.37 )
@@
$e^{-2\phi}={\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}.$ (3.38)3.38( 3.38 )
We in particular note that there is no cross term in the metric.
@@
$\gamma={\rho_{-}-\rho_{+}\over\rho_{-}+\rho_{+}},\quad\rho_{+}>\rho_{-}.$ (3.39)3.39( 3.39 )
@@
$\eqalign{d{\widehat{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{\rho_{+}^{2}\over l^{2}(1-\gamma)^{2}}\left(1-{\rho_{+}^{2}\over\rho^{2}}\right)d{\widehat{t}}^{2}+{\rho_{+}^{2}\over(1-\gamma)^{2}}\left(1-{\rho_{-}^{2}\over\rho^{2}}\right)d{\widehat{x}}^{2},\cr B&=B_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}}\wedge d{\widehat{x}},\quad B_{{\widehat{t}}{\widehat{x}}}={\rho^{2}_{+}\over l(1-\gamma)^{2}}\left({\rho_{-}\rho_{+}\over\rho^{2}}-1\right),\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|={g_{s}^{2}\over(1-\gamma)^{2}}{\rho_{+}^{2}\over\rho^{2}}.}$ (3.40)3.40( 3.40 )
These particular non-rotating solutions are already known in the literature and they are first obtained in [21]. They are exact solutions of classical string theory. This family of solutions is related to (3.24) by the Lorentz boost (3.35). To put them into the form given in [21] we introduce the new coordinates
${\rho_{+}\over l(1-\gamma)}{\widehat{t}}={l\over 2}\tau,\quad{\rho_{+}\over(1-\gamma)}{\widehat{x}}={l\over 2}\chi,\quad\rho^{2}=r.$ (3.41)3.41( 3.41 )
@@
$\eqalign{d{\overline{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over r^{2}\left(1-{r_{+}\over r}\right)\left(1-{r_{-}\over r}\right)}-\left(1-{r_{+}\over r}\right)d{\tau}^{2}+\left(1-{r_{-}\over r}\right)d{\chi}^{2}\right],\cr B&=B_{{\tau}{\chi}}d{\tau}\wedge d{\chi},\quad B_{{\tau}{\chi}}={l^{2}\over 4}\left({\sqrt{r_{-}r_{+}}\over r}-1\right),\cr-2\Phi&=\ln\left({|r|\over r_{+}}\right)+{\tenrm const.},}$ (3.42)3.42( 3.42 )
where $r_{\pm}=\rho_{\pm}^{2}$.18 The constant is ${\tenrm const.}=\ln\left[(1-\gamma)^{2}/g_{s}^{2}\right]$. It is invariant under $\gamma\to 1/\gamma$. Note $\Phi_{\tenrm there}=-2\Phi_{\tenrm here}$. Note also in general $\chi$ need not be periodic (with a definite period) and $r$ can take negative values. However here we will assume, for simplicity, $\chi$ is compact and $r$ is positive. The Ricci scalar is singular at $r=0$ and it is finite at $r=r_{\pm}$. We note that inside the event horizon, i.e., in the region $0<r<r_{+}$, the coordinate $\tau$ is everywhere space-like. We also note that $\chi$ becomes time-like in the region $0<r<r_{-}$. We will not discuss (3.42) further here. We refer to [21] for more details and a comprehensive discussion.
We first note that for $\gamma\neq 0$ the $B$ field, i.e. its component $B_{{\widehat{t}}{\widehat{x}}}$, is a constant at $\rho=\infty$. In general, we can use the gauge ambiguity in defining $B$ to set the constant at infinity to any arbitrary value. An equivalent Kalb-Ramond field, i.e., that gives the same $H=dB$, is
$\eqalign{{\widetilde{B}}_{{\widehat{t}}{\widehat{x}}}&=B_{{\widehat{t}}{\widehat{x}}}+{\lambda(\rho_{+}^{2}-\rho_{-}^{2})\over 2l},\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\gamma)+\rho_{-}^{2}(1-\gamma)-2\rho^{2}\right]+{\lambda(\rho_{+}^{2}-\rho_{-}^{2})\over 2l},\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\gamma)(1+\lambda(1+\gamma))+\rho_{-}^{2}(1-\gamma)(1-\lambda(1-\gamma))-2\rho^{2}(1+2\lambda\gamma)\right],\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\lambda+\gamma(1+2\lambda)+\lambda\gamma^{2})+\rho_{-}^{2}(1-\lambda-\gamma(1-2\lambda)-\lambda\gamma^{2})-2\rho^{2}(1+2\lambda\gamma)\right].}$ (3.43)3.43( 3.43 )
@@
We next periodically identify ${\widehat{x}}$ and perform $T$-duality. This gives the dilaton black hole solution described by the metric and dilaton,
@@
$\eqalign{d{\widetilde{s}}^{2}&=g_{\rho\rho}d\rho^{2}+{1\over g_{{\widehat{x}}{\widehat{x}}}}(d{\widetilde{x}}-{\widetilde{B}}_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}})^{2}+g_{{\widehat{t}}{\widehat{t}}}d{\widehat{t}}^{2},\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})\over l^{2}\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}^{2}\cr&+\left[{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{-}^{2})}\right]\cr&\cdot\left\{d{\widetilde{x}}-{(\rho_{+}^{2}-\rho_{-}^{2})\left[\rho_{+}^{2}(1+\gamma)(1+\lambda(1+\gamma))+\rho_{-}^{2}(1-\gamma)(1-\lambda(1-\gamma))-2\rho^{2}(1+2\lambda\gamma)\right]\over 2l\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}\right\}^{2},\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(4\lambda(1+\gamma\lambda)\rho^{2}+\rho_{-}^{2}(1-(1-\gamma)\lambda)^{2}-\rho_{+}^{2}(1+(1+\gamma)\lambda)^{2})\over 4l^{2}(\rho^{2}-\rho_{-}^{2})}d{\widehat{t}}^{2}\cr&-2{\rho_{-}^{2}(1-\gamma)(1-(1-\gamma)\lambda)+\rho_{+}^{2}(1+\gamma)(1+(1+\gamma)\lambda)-2(1+2\gamma\lambda)\rho^{2}\over 2l(\rho^{2}-\rho_{-}^{2})}d{\widetilde{x}}d{\widehat{t}}\cr&+{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{-}^{2})}d{\widetilde{x}}^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}(\rho^{2}-\rho_{-}^{2})|,}$ (3.44)3.44( 3.44 )
@@
where ${\widetilde{x}}$ is the dual coordinate. Under $\gamma\to 1/\gamma$, ${\widetilde{x}}\to-{\widetilde{x}}/\gamma$ and ${\widehat{t}}\to{\widehat{t}}/\gamma$. We note that the $\widetilde{B}$ field, and thus in turn the gauge parameter $\lambda$, enters into the coefficient of the cross term $d{\widehat{t}}d{\widetilde{x}}$. The solution is invariant under $\gamma\to 1/\gamma$ because $\lambda$ transforms as $\lambda\to-\gamma(1+\gamma\lambda)$. We will assume $0<\lambda<-1/\gamma$ so that the solution describes a black hole. The scalar curvature is given by
$R={4(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}(\rho^{2}-\rho_{-}^{2})}.$ (3.45)3.45( 3.45 )
@@
Thus, the black hole solutions (labelled by the triplet $(\rho_{-},\rho_{+};\lambda)$) have a curvature singularity enclosed inside their event horizons.
@@
We specially note that $R$ does not depend on the coupling $\gamma$. In particular for $\lambda=-1/(2\gamma)$ and $\gamma=(\rho_{-}-\rho_{+})/(\rho_{-}+\rho_{+})$ the solution can be put using the coordinates (3.41) into the form
@@
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over\left(1-{r_{+}\over r}\right)\left(1-{r_{-}\over r}\right)r^{2}}-\left(1-{r_{+}\over r}\right)d\tau^{2}+{1\over\left(1-{r_{-}\over r}\right)}\left(d\chi-{\sqrt{r_{-}r_{+}}\over r}d\tau\right)^{2}\right],\cr-2\widetilde{\Phi}&=\ln(|r-r_{-}|)+{\tenrm const.},}$ (3.46)3.46( 3.46 )
where $r_{\pm}=\rho_{\pm}^{2}$. In the limit $r_{+}\to r_{-}:=r_{0}$ this simplifies to
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over\left(1-{r_{0}\over r}\right)^{2}r^{2}}-\left(1-{r_{0}\over r}\right)d\tau^{2}+{1\over\left(1-{r_{0}\over r}\right)}\left(d\chi-{r_{0}\over r}d\tau\right)^{2}\right],\cr-2\widetilde{\Phi}&=\ln(|r-r_{0}|)+{\tenrm const.}.}$ (3.47)3.47( 3.47 )
@@
We note from (3.45) that in this limit the string frame Ricci scalar $R=0$. However, we also note that the dilaton diverges at $r=r_{0}$. In fact, in the Einstein frame, the scalar curvature depends on $r$ and it is singular at $r=r_{0}$. It is given by $R=-8/(l^{2}(r-r_{0})^{2})$. Thus, (in the Einstein frame) the solution (3.47) describes a naked singularity. In this paper we in general assume $\rho_{+}>\rho_{-}$.

### v5

@@ 3.3. Positive coupling
Fig. 4: The plot depicts the global structure of the black string solutions with $10<\gamma<$ 1. $\rho_{\pm}$ represent the horizons. $\rho=0$ is a causal singularity. It is similar to the causal singularity in BTZ black hole. The solutions have a time-like singularity at $\rho=\rho_{0}$. $\rho_{0}$ is represented by zigzag lines. For arbitrarily small positive $\gamma$, the singularity is at arbitrarily large $\rho=\rho_{0}$.
@@
$\eqalign{ds^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho_{+}^{2}-\rho^{2})(\rho_{-}^{2}-\rho^{2})}-{(\rho^{2}_{+}-\rho^{2}_{-})(\rho^{2}_{-}-\rho^{2})dt^{2}\over 4l^{2}\rho^{2}}+{(\rho_{+}^{2}-\rho_{-}^{2})\rho^{2}\over 4(\rho^{2}_{+}-\rho^{2})}\left(dx-{\rho_{-}\rho_{+}\over l\rho^{2}}dt\right)^{2},\cr B&=B_{tx}dt\wedge dx,\quad B_{tx}={\rho_{+}^{2}-\rho_{-}^{2}\over 4l}={\tenrm const.},\quad Q=0,\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,\quad e^{-2\phi}=4\left({\rho_{+}^{2}-\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right).}$ (3.40)3.40( 3.40 )
@@
$R={4\over l^{2}}\left({\rho_{+}^{2}-\rho_{-}^{2}\over\rho^{2}-\rho_{+}^{2}}\right).$ (3.41)3.41( 3.41 )
Thus, for an observer residing in the spacetime region $\rho<\rho_{-}$, i.e., the exterior region, the solution (3.40) has a singularity at $\rho=\rho_{0}=\rho_{+}$ beyond the horizon, i.e., $\rho_{0}>\rho_{-}$. Thus, in view of this observer, (3.40) describes a black hole. Since $x$ is compact, the solution contains closed time-likes curve in the region $\rho>\rho_{0}=\rho_{+}$. Thus, they are located in the interior region. We note that the $B$ field, i.e. $B_{tx}$, is a non-zero constant. We can fix the $B$ field to zero using the gauge freedom. The Penrose diagram is given in Fig. 5.
Fig. 5: The plot depicts the global structure of the black string solution with $\gamma=1$. $\rho_{-}$ represents the horizon. $\rho=0$ is a causal singularity. The solution has a space-like singularity at $\rho_{0}=\rho_{+}$. $\rho_{0}$ is represented by zigzag lines.
@@ 3.4. New black hole solutions
$\eqalign{t&=\widehat{t}{\tenrm cosh}(\alpha)+l\widehat{x}{\tenrm sinh}(\alpha),\cr lx&=l\widehat{x}{\tenrm cosh}(\alpha)+\widehat{t}{\tenrm sinh}(\alpha),}$ (3.42)3.42( 3.42 )
@@
$\tanh(2\alpha)={2\rho_{+}\rho_{-}\over\rho_{+}^{2}+\rho_{-}^{2}}.$ (3.43)3.43( 3.43 )
@@
$\eqalign{d{\widehat{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-l^{-2}e^{2\phi}(\rho^{2}-\rho_{+}^{2})d{\widehat{t}}^{2}+e^{2\phi}(\rho^{2}-\rho_{-}^{2})d{\widehat{x}}^{2},\cr B&=B_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}}\wedge d{\widehat{x}},\cr B_{{\widehat{t}}{\widehat{x}}}&={1\over 2}e^{2\phi}\left({\rho_{+}^{2}-\rho_{-}^{2}\over l}\right)\left[\gamma-1-2\left({\rho^{2}-\rho_{+}^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}\right)\right],\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\gamma)+\rho_{-}^{2}(1-\gamma)-2\rho^{2}\right],\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|,}$ (3.44)3.44( 3.44 )
@@
$e^{-2\phi}={\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over\rho_{+}^{2}-\rho_{-}^{2}}.$ (3.45)3.45( 3.45 )
We in particular note that there is no cross term in the metric. Therefore, in the case $\widehat{x}$ is periodic they describe non-rotating black strings.22 At $\gamma=0$, the dilaton is constant and the metric is (equivalent to) the non-rotating BTZ black hole. $\rho=\rho_{-}$ is a causal singularity. That is, in the case ${\widehat{x}}$ is periodic, there are closed time-like curves in the region $0\leq\rho<\rho_{-}$. At $\gamma=-1$, the coefficient of $d{\widehat{x}}^{2}$ equals $\left(\rho^{2}_{+}-\rho^{2}_{-}\right)/4$, i.e., it is a constant. One may as well prefer to refer to the solutions (3.44) as deformed black holes.
@@
$\gamma={\rho_{-}-\rho_{+}\over\rho_{-}+\rho_{+}},\quad\rho_{+}>\rho_{-}.$ (3.46)3.46( 3.46 )
@@
$\eqalign{d{\widehat{s}}^{2}&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{\rho_{+}^{2}\over l^{2}(1-\gamma)^{2}}\left(1-{\rho_{+}^{2}\over\rho^{2}}\right)d{\widehat{t}}^{2}+{\rho_{+}^{2}\over(1-\gamma)^{2}}\left(1-{\rho_{-}^{2}\over\rho^{2}}\right)d{\widehat{x}}^{2},\cr B&=B_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}}\wedge d{\widehat{x}},\quad B_{{\widehat{t}}{\widehat{x}}}={\rho^{2}_{+}\over l(1-\gamma)^{2}}\left({\rho_{-}\rho_{+}\over\rho^{2}}-1\right),\cr e^{2\Phi}&=g_{s}^{2}|e^{2\phi}|={g_{s}^{2}\over(1-\gamma)^{2}}{\rho_{+}^{2}\over\rho^{2}}.}$ (3.47)3.47( 3.47 )
These particular non-rotating solutions (with independent $\rho_{\pm}$ parameters) are already known in the literature and they are first obtained in [23]. They are exact solutions of classical string theory. This family of solutions is related to (3.31), thus also to (3.23), by the Lorentz boost (3.42). Note $\gamma$ takes the special value (3.46), see also (3.30) and (3.26). To put them into the form given in [23] we introduce the new coordinates
${\rho_{+}\over l(1-\gamma)}{\widehat{t}}={l\over 2}\tau,\quad{\rho_{+}\over(1-\gamma)}{\widehat{x}}={l\over 2}\chi,\quad\rho^{2}=r.$ (3.48)3.48( 3.48 )
@@
$\eqalign{d{\overline{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over r^{2}\left(1-{r_{+}\over r}\right)\left(1-{r_{-}\over r}\right)}-\left(1-{r_{+}\over r}\right)d{\tau}^{2}+\left(1-{r_{-}\over r}\right)d{\chi}^{2}\right],\cr B&=B_{{\tau}{\chi}}d{\tau}\wedge d{\chi},\quad B_{{\tau}{\chi}}={l^{2}\over 4}\left({\sqrt{r_{-}r_{+}}\over r}-1\right),\cr-2\Phi&=\ln\left({|r|\over r_{+}}\right)+{\tenrm const.},\quad{r_{-}\over r_{+}}=\left({1+\gamma\over 1-\gamma}\right)^{2},}$ (3.49)3.49( 3.49 )
where $r_{\pm}=\rho_{\pm}^{2}$. We could more generally take $r_{-}$ and $r_{+}$ as independent parameters. That is, we could simply disregard the constraint on the ratio $r_{-}/r_{+}$, see also (3.46). However, this would correspond to a (fundamentally) different case. Note $\Phi_{\tenrm there}=-2\Phi_{\tenrm here}$.23 The constant is ${\tenrm const.}=\ln\left[(1-\gamma)^{2}/g_{s}^{2}\right]$. It is invariant under $\gamma\to 1/\gamma$. Note also in general $\chi$ need not be periodic (with a definite period) and $r$ can take negative values. However here we will assume, for simplicity, $\chi$ is compact and $r$ is positive. The Ricci scalar $R$ is singular at $r=0$ and it is finite at $r=r_{\pm}$. It is given by $R=2(2r(r_{-}+r_{+})-7r_{-}r_{+})/(lr)^{2}$. We note that inside the event horizon, i.e., in the region $0<r<r_{+}$, the coordinate $\tau$ is everywhere space-like. We also note that $\chi$ becomes time-like in the region $0<r<r_{-}$. We will not discuss (3.49) further here. We refer to [23] for more details and a comprehensive discussion.
We first note that for $\gamma\neq 0$ the $B$ field, i.e. its component $B_{{\widehat{t}}{\widehat{x}}}$ (3.44), is a constant at $\rho=\infty$. In general, we can use the gauge ambiguity in defining $B$ to set the constant at infinity to any arbitrary value. An equivalent Kalb-Ramond field, i.e., that gives the same $H=dB$, is
$\eqalign{{\widetilde{B}}_{{\widehat{t}}{\widehat{x}}}&=B_{{\widehat{t}}{\widehat{x}}}+{\lambda(\rho_{+}^{2}-\rho_{-}^{2})\over 2l},\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\gamma)+\rho_{-}^{2}(1-\gamma)-2\rho^{2}\right]+{\lambda(\rho_{+}^{2}-\rho_{-}^{2})\over 2l},\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\gamma)(1+\lambda(1+\gamma))+\rho_{-}^{2}(1-\gamma)(1-\lambda(1-\gamma))-2\rho^{2}(1+2\lambda\gamma)\right],\cr&={1\over 2l}e^{2\phi}\left[\rho_{+}^{2}(1+\lambda+\gamma(1+2\lambda)+\lambda\gamma^{2})+\rho_{-}^{2}(1-\lambda-\gamma(1-2\lambda)-\lambda\gamma^{2})-2\rho^{2}(1+2\lambda\gamma)\right].}$ (3.50)3.50( 3.50 )
@@
We next periodically identify ${\widehat{x}}$ and perform $T$-duality. This gives (in units where $\alpha^{\prime}=1$) the dilaton black hole solution described by the metric and dilaton,
@@
$\eqalign{d{\widetilde{s}}^{2}&=g_{\rho\rho}d\rho^{2}+{1\over g_{{\widehat{x}}{\widehat{x}}}}(d{\widetilde{x}}-{\widetilde{B}}_{{\widehat{t}}{\widehat{x}}}d{\widehat{t}})^{2}+g_{{\widehat{t}}{\widehat{t}}}d{\widehat{t}}^{2},\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{+}^{2})\over l^{2}\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}^{2}\cr&+\left[{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{-}^{2})}\right]\cr&\cdot\left\{d{\widetilde{x}}-{(\rho_{+}^{2}-\rho_{-}^{2})\left[\rho_{+}^{2}(1+\gamma)(1+\lambda(1+\gamma))+\rho_{-}^{2}(1-\gamma)(1-\lambda(1-\gamma))-2\rho^{2}(1+2\lambda\gamma)\right]\over 2l\left[\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\right]}d{\widehat{t}}\right\}^{2},\cr&={l^{2}\rho^{2}d\rho^{2}\over(\rho^{2}-\rho_{+}^{2})(\rho^{2}-\rho_{-}^{2})}-{(\rho_{+}^{2}-\rho_{-}^{2})(4\lambda(1+\gamma\lambda)\rho^{2}+\rho_{-}^{2}(1-(1-\gamma)\lambda)^{2}-\rho_{+}^{2}(1+(1+\gamma)\lambda)^{2})\over 4l^{2}(\rho^{2}-\rho_{-}^{2})}d{\widehat{t}}^{2}\cr&-2{\rho_{-}^{2}(1-\gamma)(1-(1-\gamma)\lambda)+\rho_{+}^{2}(1+\gamma)(1+(1+\gamma)\lambda)-2(1+2\gamma\lambda)\rho^{2}\over 2l(\rho^{2}-\rho_{-}^{2})}d{\widetilde{x}}d{\widehat{t}}\cr&+{\rho_{+}^{2}(1+\gamma)^{2}-\rho_{-}^{2}(1-\gamma)^{2}-4\gamma\rho^{2}\over(\rho_{+}^{2}-\rho_{-}^{2})(\rho^{2}-\rho_{-}^{2})}d{\widetilde{x}}^{2},\cr e^{-2\widetilde{\Phi}}&=|g_{s}^{-2}(\rho^{2}-\rho_{-}^{2})|,}$ (3.51)3.51( 3.51 )
@@
where ${\widetilde{x}}$ is the dual coordinate. Under $\gamma\to 1/\gamma$, ${\widetilde{x}}\to-{\widetilde{x}}/\gamma$ and ${\widehat{t}}\to{\widehat{t}}/\gamma$. We note that the $\widetilde{B}$ field, and thus in turn the gauge parameter $\lambda$, enters into the coefficient of the cross term $d{\widehat{t}}d{\widetilde{x}}$. The solution is invariant under $\gamma\to 1/\gamma$ because $\lambda$ transforms as $\lambda\to-\gamma(1+\gamma\lambda)$. We will assume $0<\lambda<-1/\gamma$ so that the solution describes a black hole. That is, we require the sign of the time component of the metric, i.e. the coefficient of $d{\widehat{t}}^{2}$, to be negative also on the section ${\widetilde{x}}={\tenrm constant}$ in the region $\rho>\rho_{+}$. The scalar curvature is given by
$R={4(\rho_{+}^{2}-\rho_{-}^{2})\over l^{2}(\rho^{2}-\rho_{-}^{2})}.$ (3.52)3.52( 3.52 )
@@
Thus, the black hole solutions (labelled by the quartet $(\rho_{-},\rho_{+};\gamma;\lambda)$) have a curvature singularity enclosed inside their event horizons.
@@
We specially note that $R$ does not depend on the coupling $\gamma$. In particular for $\lambda=-1/(2\gamma)$ and, $\gamma=(\rho_{-}-\rho_{+})/(\rho_{-}+\rho_{+})$ the solution can be put using the coordinates (3.48) into the form
@@
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over\left(1-{r_{+}\over r}\right)\left(1-{r_{-}\over r}\right)r^{2}}-\left(1-{r_{+}\over r}\right)d\tau^{2}+{1\over\left(1-{r_{-}\over r}\right)}\left(d\chi-{\sqrt{r_{-}r_{+}}\over r}d\tau\right)^{2}\right],\cr-2\widetilde{\Phi}&=\ln(|r-r_{-}|)+{\tenrm const.},}$ (3.53)3.53( 3.53 )
where $r_{\pm}=\rho_{\pm}^{2}$ and ${\tenrm const.}=-2\ln g_{s}$. In the (extremal) limit $r_{+}\to r_{-}:=r_{0}$ this simplifies to
$\eqalign{d{\widetilde{s}}^{2}&={l^{2}\over 4}\left[{dr^{2}\over\left(1-{r_{0}\over r}\right)^{2}r^{2}}-\left(1-{r_{0}\over r}\right)d\tau^{2}+{1\over\left(1-{r_{0}\over r}\right)}\left(d\chi-{r_{0}\over r}d\tau\right)^{2}\right],\cr-2\widetilde{\Phi}&=\ln(|r-r_{0}|)+{\tenrm const.}.}$ (3.54)3.54( 3.54 )
@@
We note from (3.52) that in this limit the string frame Ricci scalar $R=0$. The Kretschmann scalar can also be shown to be zero. However, we also note that the dilaton diverges at $r=r_{0}$. Therefore, near $r=r_{0}$ the string coupling is large. In fact, in the Einstein frame, the scalar curvature depends on $r$ and it is singular at $r=r_{0}$. It is given by $R=-8g_{s}^{4}/(l^{2}(r-r_{0})^{2})$. Thus, in the Einstein frame, the solution (3.54) describes a naked singularity.24 The solution in the Einstein frame is $\eqalign{d{\widetilde{s}}^{2}&={1\over 4}\cdot{k\over g_{s}^{4}}\cdot\left[dr^{2}-r^{2}\left(1-{r_{0}\over r}\right)^{3}d\tau^{2}+r^{2}\left(1-{r_{0}\over r}\right)\left(d\chi-{r_{0}\over r}d\tau\right)^{2}\right],\cr-2\widetilde{\Phi}&=\ln(|r-r_{0}|)-2\ln g_{s},\quad g_{s}^{2}={8\pi G_{N}\over\kappa_{0}^{2}},\quad k={l^{2}\over l_{s}^{2}},\quad l_{s}\propto l_{p}.}$ (3.55)3.55( 3.55 ) $l_{p}$ is Planck length. The geometry that the string metric (3.54) describes, for $r>r_{0}$ and constant $\tau$, looks like a funnel. Thus, the strings get stretched more and more as $r$ gets closer and closer to $r_{0}$. Note in general $\chi$ need not be periodic with a definite period. We leave a further study of the (intriguing) solution (3.54) to a future work. In this paper we in general assume $\rho_{+}>\rho_{-}$.